# Stable High-Order Interpolation on the Grassmann Manifold by Maximum-Volume Coordinates and Arnoldi Orthogonalization

### 25 Mar 2026 v1

@@ 2 Interpolation on the Grassmann Manifold
2.1 Problem Formulation
Let $f:D\rightarrow\operatorname{Gr}(n,p)$ be a differentiable manifold-valued function defined on a parameter domain $D\subset\mathbb{R}$. Given a discrete data set $\{(t_{i},\mathcal{U}_{i})\}_{i=1}^{m}$ sampled from the function, where $t_{i}\in\mathbb{R}$ are parameter nodes and $\mathcal{U}_{i}\in\operatorname{Gr}(n,p)$ are the corresponding subspaces, the goal is to construct an interpolant $\hat{f}:D\rightarrow\operatorname{Gr}(n,p)$ such that $\hat{f}(t_{i})=\mathcal{U}_{i}$ for all $i$.
@@ 2.1 Problem Formulation
2.2 Maximum-Volume Coordinates: A Matrix-Free Geometric Mapping
The maximum-volume (MV) coordinate [19] is a matrix-decomposition-free alternative to Riemannian normal coordinates for representing subspaces on the Grassmann manifold. The approach exploits the natural block structure of Stiefel representatives to construct local coordinates using only algebraic operations, and avoids the singular value decompositions.
2.2.1 Local Coordinate Representation
Let $\operatorname{Gr}(n,p)$ denote the Grassmann manifold of $p$-dimensional subspaces in $\mathbb{R}^{n}$, represented numerically through Stiefel matrices $U\in\operatorname{St}(n,p)$ with orthonormal columns. Each subspace $\mathcal{U}\in\operatorname{Gr}(n,p)$ corresponds to an equivalence class $U]=\{UQ:Q\in O(p)\}$. Given a sample basis $U\in\operatorname{St}(n,p)$ and a row permutation $P$, we partition
@@ 2.2.3 Derivative Mapping for Hermite Data
Applying the chain rule to the MV representation $\Xi(t)=\widetilde{U}_{2}(t)\widetilde{U}_{1}(t)^{-1}$ and evaluating at $t=0$, the MV-coordinate derivative satisfies [19, 25]:
@@
2.2.4 Inverse Mapping and Reconstruction
Given an interpolated coordinate matrix $\Xi$, a Stiefel representative can be reconstructed using the inverse of the MV-representation. From (1), we have $\widetilde{U}_{2}=\Xi\widetilde{U}_{1}$. Combined with the orthonormality condition $\widetilde{U}_{1}^{T}\widetilde{U}_{1}+\widetilde{U}_{2}^{T}\widetilde{U}_{2}=I_{p}$, we obtain $\widetilde{U}_{1}^{T}(I_{p}+\Xi^{T}\Xi)\widetilde{U}_{1}=I_{p}$. A numerically stable reconstruction is then:
@@ 2.2.4 Inverse Mapping and Reconstruction
To maintain computational efficiency during reconstruction, we avoid explicitly computing the symmetric inverse square root $(I_{p}+\Xi^{T}\Xi)^{-1/2}$, which necessitates eigenvalue decompositions. Instead, we utilize the Cholesky factorization $LL^{T}=I_{p}+\Xi(s)^{T}\Xi(s)$. By defining the inverse factor $M=L^{-T}$, the reconstructed Stiefel representative is given by:
@@ 3.1 The V+A Method
where $\star$ denotes the scalar $h_{n+1,n}e_{n}^{T}H_{n}^{n-1}e_{1}$. In matrix form, the Arnoldi process thus yields a QR-style factorization of the Vandermonde matrix:
@@
where $R_{n+1}\in\mathbb{R}^{(n+1)\times(n+1)}$ is an implicitly defined upper triangular matrix:
@@
Given the data fitting problem $V_{x}\mathbf{c}=\mathbf{f}$ (or $V_{x}\mathbf{c}\approx\mathbf{f}$ in the least-squares sense), substituting the factorization (5) yields $Q_{n+1}R_{n+1}\mathbf{c}=\mathbf{f}$. By defining the new coefficient vector $\mathbf{d}=R_{n+1}\mathbf{c}$, the system is transformed from an ill-conditioned monomial basis into a well conditioned orthogonal basis. Since $Q_{n+1}$ has orthonormal columns, $\mathbf{d}$ is computed reliably via orthogonal projection:
@@
To evaluate the interpolant at new points $s_{1},\ldots,s_{M}$, we construct an evaluation basis $W_{n+1}\in\mathbb{R}^{M\times(n+1)}$ that obeys the exact recurrence relation encoded in $\tilde{H}_{n}$. Let $S=\operatorname{diag}(s_{1},\ldots,s_{M})$. Starting with $w_{1}=\mathbf{e}_{M}/\sqrt{M}$, the subsequent columns for $k=1,\ldots,n$ are generated via:
@@
This ensures that $W_{n+1}$ spans the Krylov subspace $\mathcal{K}_{n+1}(S,\mathbf{e}_{M})$ and satisfies the relation:
@@ 3.2 The Confluent V+A Method
Executing the Arnoldi process for $n$ steps generates an orthonormal basis $\mathcal{Q}_{n+1}\in\mathbb{R}^{2m\times(n+1)}$ and an upper Hessenberg matrix $\mathcal{H}_{n}\in\mathbb{R}^{(n+1)\times n}$ governed by the relation:
@@
where $R_{n+1}$ is an upper triangular matrix that is never explicitly formed. Due to the block structure of the augmented space, the orthonormal basis $\mathcal{Q}_{n+1}$ naturally partitions into function and derivative domains:
@@
For the augmented data fitting problem $\begin{bmatrix}V\\ V_{(1)}\end{bmatrix}\mathbf{c}\approx\begin{bmatrix}\mathbf{f}\\ \dot{\mathbf{f}}\end{bmatrix}$, substituting (8) shifts the problem into the well-conditioned orthogonal basis. We define the stable coefficient vector $\mathbf{d}=R_{n+1}\mathbf{c}$, which is efficiently computed via orthogonal projection:
@@
For subspace-valued manifold data mapped to the Euclidean coordinate domain, we consider the block matrix $\mathbf{Y}=\begin{pmatrix}\mathbf{X}\\ \dot{\mathbf{X}}\end{pmatrix}\in\mathbb{R}^{2m\times N_{\text{coord}}}$, where the rows of $\mathbf{X}$ and $\dot{\mathbf{X}}$ contain the vectorized MV coordinates $\mathbf{x}_{i}^{T}$ and their velocities $\dot{\mathbf{x}}_{i}^{T}$. The corresponding coefficient matrix is computed analogously:
@@
To evaluate the Hermite interpolant at new points $s_{1},\ldots,s_{M}$, we construct an evaluation basis $\mathcal{W}_{n+1}\in\mathbb{R}^{2M\times(n+1)}$. Let $S=\operatorname{diag}(s_{1},\ldots,s_{M})$ and define the augmented evaluation operator:
@@
This recurrence mirrors the original Arnoldi process, ensuring that $\mathcal{W}_{n+1}$ spans the augmented Krylov subspace $\mathcal{K}_{n+1}(\mathcal{S},\mathcal{W}_{1})$ and satisfies:
@@
Consequently, the evaluation basis also partitions into function and derivative components:
@@
In this architecture, $\mathcal{H}_{n}$ acts as an algebraic bridge, transporting the coupled derivative recurrence structures from the sample nodes to the evaluation points. The interpolated manifold coordinates and their velocities are finally obtained via:
@@ 4.1 Stage I: Geometric Stabilization via Orthogonal Transformation
  • Coordinate extraction. All samples are transformed by the same orthogonal map:
@@ 4.2 Stage II: Algebraic Interpolation via Krylov Subspaces
The Vandermonde with Arnoldi (V+A) method circumvents the problem by replacing the monomial basis with an orthogonal basis generated from the nodes. The column space of $V$ is a Krylov subspace. With $D=\operatorname{diag}(t_{1},\dots,t_{m})$, the Arnoldi process produces:
@@
Coefficient extraction reduces to orthogonal projection:
@@ 4.3 Stage III: Geometric Reconstruction
  • Inverse transformation. The global orthogonal map $\mathcal{Q}$ (from Stage I) is applied to return the subspace to its original orientation:
@@ 5 Convergence and Error Analysis
Meanwhile, the derivative of the right-hand side is $d(I_{p}+\Xi^{T}\Xi)|_{0}=\Delta^{T}\Xi+\Xi^{T}\Delta\big|_{\Xi=0}=0$. Consequently, $2dS|_{0}=0$, which implies $dS|_{0}=0$. By applying the product rule to the reconstruction formula $\widetilde{U}(\Xi)=\begin{bmatrix}I_{p}\\ \Xi\end{bmatrix}S(\Xi)$ (see Lemma 1), the differential in direction $\Delta$ is:
@@ 5.1 Error Propagation in Coordinate Systems
Let $\Xi$ be the true MV coordinates and $\hat{\Xi}=\Xi+\Delta$ be the interpolated coordinates. Under the Householder-stabilized system, the reconstruction error for the orthogonal projector $P=UU^{T}$ satisfies:
@@
Let $\Psi(\Xi)=\widetilde{U}(\Xi)\widetilde{U}(\Xi)^{T}$ be the mapping from the coordinate domain to the Grassmannian. Define the line segment $\gamma(t)=\Xi+t\Delta$ for $t\in[0,1]$. By the Mean Value Theorem for matrix-valued functions:
@@
Taking the Frobenius norm on both sides yields:
@@
To bound the differential norm at any point $X\in\{\gamma(t)\}$, we evaluate $d\Psi(X)[\Delta]=(d\widetilde{U})\widetilde{U}^{T}+\widetilde{U}(d\widetilde{U})^{T}$. Using the reconstruction formula $\widetilde{U}(X)=\begin{bmatrix}I_{p}\\ X\end{bmatrix}S(X)$ where $S(X)=(I_{p}+X^{T}X)^{-1/2}$, the differential is:
@@
Its unique solution is given by the integral representation:
@@
Taking the spectral norm and applying the bound $|e^{-S^{-1}\tau}|_{2}\leq e^{-\tau/|S|_{2}}$, the scalar integral can be evaluated to produce:
@@
Substituting $d\tilde{U}$ into $d\Psi$, we decompose it into a principal part and a remainder:
@@
Let $A=D_{1}\tilde{U}^{T}=\begin{bmatrix}0\\ \Delta S\end{bmatrix}\tilde{U}^{T}$. Applying the triangle inequality to the remainder $\mathcal{R}=D_{2}\tilde{U}^{T}+\tilde{U}D_{2}^{T}$ gives:
@@
Thus the remainder is rigorously bounded by $\|\mathcal{R}\|_{F}=\mathcal{O}(\|X\|_{2}\|\Delta\|_{F})$. Next, we evaluate $\|A+A^{T}\|_{F}$. By the properties of the Frobenius norm:
@@
Thus, $\operatorname{tr}(A^{2})=\operatorname{tr}(SX^{T}\Delta S^{2}X^{T}\Delta S)$. By Cauchy-Schwarz inequality for trace inner products $|\operatorname{tr}(YZ)|\leq\|Y\|_{F}\|Z\|_{F}$, letting $Y=SX^{T}\Delta S$ and $Z=SX^{T}\Delta S$, we get:
@@
Substituting back into (16), we obtain:
@@
Using the sub-multiplicative property $\|\Delta S\|_{F}\leq\|S\|_{2}\|\Delta\|_{F}$:
@@
Combining the bounds for the principal part and the remainder $\mathcal{R}$, the differential norm at $X$ is bounded by:
@@
Returning to the supremum in (14), since $X\in\{\Xi+t\Delta\}$, we have $\|X\|_{2}\leq\|\Xi\|_{2}+\|\Delta\|_{2}$. Furthermore, as established in Lemma 2, the matrix function $S(\cdot)$ is smooth, ensuring $\|S(X)\|_{2}=\|S(\Xi)\|_{2}+\mathcal{O}(\|\Delta\|_{F})$. Since the upper block of $\widetilde{U}(\Xi)$ is $S(\Xi)$, we have $\|\widetilde{U}_{1}\|_{2}=\|S(\Xi)\|_{2}$. Substituting these limits into the integral bound completes the strict derivation:

### 27 Mar 2026 v2

@@ 2 Interpolation on the Grassmann Manifold
Let $\operatorname{Gr}(n,p)$ denote the Grassmann manifold of $p$-dimensional subspaces in $\mathbb{R}^{n}$, represented numerically through Stiefel matrices $U\in\operatorname{St}(n,p)$ with orthonormal columns. Each subspace $\mathcal{U}\in\operatorname{Gr}(n,p)$ corresponds to an equivalence class $U]=\{UQ:Q\in O(p)\}$. Let $f:D\rightarrow\operatorname{Gr}(n,p)$ be a differentiable manifold-valued function defined on a parameter domain $D\subset\mathbb{R}$. Given a discrete data set $\{(t_{i},\mathcal{U}_{i})\}_{i=1}^{m}$ sampled from the function, where $t_{i}\in\mathbb{R}$ are parameter nodes and $\mathcal{U}_{i}\in\operatorname{Gr}(n,p)$ are the corresponding subspaces, the goal is to construct an interpolant $\hat{f}:D\rightarrow\operatorname{Gr}(n,p)$ such that $\hat{f}(t_{i})=\mathcal{U}_{i}$ for all $i$.
@@
The maximum-volume (MV) coordinate [19] is a matrix-decomposition-free alternative to Riemannian normal coordinates for representing subspaces on the Grassmann manifold. The approach exploits the natural block structure of Stiefel representatives to construct local coordinates using only algebraic operations, and avoids the singular value decompositions. Given a sample basis $U\in\operatorname{St}(n,p)$ and a row permutation $P$, we partition
@@
Applying the chain rule to the MV representation $\Xi(t)=\widetilde{U}_{2}(t)\widetilde{U}_{1}(t)^{-1}$ and evaluating at $t=0$, the MV-coordinate derivative satisfies [19, 25]
@@
Given an interpolated coordinate matrix $\Xi$, a Stiefel representative can be reconstructed using the inverse of the MV-representation. From (1), we have $\widetilde{U}_{2}=\Xi\widetilde{U}_{1}$. Combined with the orthonormality condition $\widetilde{U}_{1}^{T}\widetilde{U}_{1}+\widetilde{U}_{2}^{T}\widetilde{U}_{2}=I_{p}$, we obtain $\widetilde{U}_{1}^{T}(I_{p}+\Xi^{T}\Xi)\widetilde{U}_{1}=I_{p}$. A numerically stable reconstruction is then
@@
To maintain computational efficiency during reconstruction, we avoid explicitly computing the symmetric inverse square root $(I_{p}+\Xi^{T}\Xi)^{-1/2}$, which necessitates eigenvalue decompositions. Instead, we utilize the Cholesky factorization $LL^{T}=I_{p}+\Xi(s)^{T}\Xi(s)$. By defining the inverse factor $M=L^{-T}$, the reconstructed Stiefel representative is given by
@@ 3.1 The V+A Method
where $\star$ denotes the scalar $h_{n+1,n}e_{n}^{T}H_{n}^{n-1}e_{1}$. In matrix form, the Arnoldi process thus yields a QR-style factorization of the Vandermonde matrix
@@
where $R_{n+1}\in\mathbb{R}^{(n+1)\times(n+1)}$ is an implicitly defined upper triangular matrix
@@
Given the data fitting problem $V_{x}\mathbf{c}=\mathbf{f}$ (or $V_{x}\mathbf{c}\approx\mathbf{f}$ in the least-squares sense), substituting the factorization (5) yields $Q_{n+1}R_{n+1}\mathbf{c}=\mathbf{f}$. By defining the new coefficient vector $\mathbf{d}=R_{n+1}\mathbf{c}$, the system is transformed from an ill-conditioned monomial basis into a well conditioned orthogonal basis. Since $Q_{n+1}$ has orthonormal columns, $\mathbf{d}$ is computed reliably via orthogonal projection
@@
To evaluate the interpolant at new points $s_{1},\ldots,s_{M}$, we construct an evaluation basis $W_{n+1}\in\mathbb{R}^{M\times(n+1)}$ that obeys the exact recurrence relation encoded in $\tilde{H}_{n}$. Let $S=\operatorname{diag}(s_{1},\ldots,s_{M})$. Starting with $w_{1}=\mathbf{e}_{M}/\sqrt{M}$, the subsequent columns for $k=1,\ldots,n$ are generated via
@@
This ensures that $W_{n+1}$ spans the Krylov subspace $\mathcal{K}_{n+1}(S,\mathbf{e}_{M})$ and satisfies the relation
@@ 3.2 The Confluent V+A Method
Executing the Arnoldi process for $n$ steps generates an orthonormal basis $\mathcal{Q}_{n+1}\in\mathbb{R}^{2m\times(n+1)}$ and an upper Hessenberg matrix $\mathcal{H}_{n}\in\mathbb{R}^{(n+1)\times n}$ governed by the relation
@@
where $R_{n+1}$ is an upper triangular matrix that is never explicitly formed. Due to the block structure of the augmented space, the orthonormal basis $\mathcal{Q}_{n+1}$ naturally partitions into function and derivative domains
@@
For the augmented data fitting problem $\begin{bmatrix}V\\ V_{(1)}\end{bmatrix}\mathbf{c}\approx\begin{bmatrix}\mathbf{f}\\ \dot{\mathbf{f}}\end{bmatrix}$, substituting (8) shifts the problem into the well-conditioned orthogonal basis. We define the stable coefficient vector $\mathbf{d}=R_{n+1}\mathbf{c}$, which is efficiently computed via orthogonal projection
@@
For subspace-valued manifold data mapped to the Euclidean coordinate domain, we consider the block matrix $\mathbf{Y}=\begin{pmatrix}\mathbf{X}\\ \dot{\mathbf{X}}\end{pmatrix}\in\mathbb{R}^{2m\times N_{\text{coord}}}$, where the rows of $\mathbf{X}$ and $\dot{\mathbf{X}}$ contain the vectorized MV coordinates $\mathbf{x}_{i}^{T}$ and their velocities $\dot{\mathbf{x}}_{i}^{T}$. The corresponding coefficient matrix is computed analogously
@@
To evaluate the Hermite interpolant at new points $s_{1},\ldots,s_{M}$, we construct an evaluation basis $\mathcal{W}_{n+1}\in\mathbb{R}^{2M\times(n+1)}$. Let $S=\operatorname{diag}(s_{1},\ldots,s_{M})$ and define the augmented evaluation operator
@@
This recurrence mirrors the original Arnoldi process, ensuring that $\mathcal{W}_{n+1}$ spans the augmented Krylov subspace $\mathcal{K}_{n+1}(\mathcal{S},\mathcal{W}_{1})$ and satisfies
@@
Consequently, the evaluation basis also partitions into function and derivative components
@@
In this architecture, $\mathcal{H}_{n}$ acts as an algebraic bridge, transporting the coupled derivative recurrence structures from the sample nodes to the evaluation points. The interpolated manifold coordinates and their velocities are finally obtained via
@@ 4.1 Stage I: Geometric Stabilization via Orthogonal Transformation
  • Coordinate extraction. All samples are transformed by the same orthogonal map
@@ 4.2 Stage II: Algebraic Interpolation via Krylov Subspaces
The Vandermonde with Arnoldi (V+A) method circumvents the problem by replacing the monomial basis with an orthogonal basis generated from the nodes. The column space of $V$ is a Krylov subspace. With $D=\operatorname{diag}(t_{1},\dots,t_{m})$, the Arnoldi process produces
@@
Coefficient extraction reduces to orthogonal projection
@@ 4.3 Stage III: Geometric Reconstruction
  • Inverse transformation. The global orthogonal map $\mathcal{Q}$ (from Stage I) is applied to return the subspace to its original orientation
@@ 5 Convergence and Error Analysis
Meanwhile, the derivative of the right-hand side is $d(I_{p}+\Xi^{T}\Xi)|_{0}=\Delta^{T}\Xi+\Xi^{T}\Delta\big|_{\Xi=0}=0$. Consequently, $2dS|_{0}=0$, which implies $dS|_{0}=0$. By applying the product rule to the reconstruction formula $\widetilde{U}(\Xi)=\begin{bmatrix}I_{p}\\ \Xi\end{bmatrix}S(\Xi)$ (see Lemma 1), the differential in direction $\Delta$ is
@@ 5.1 Error Propagation in Coordinate Systems
Let $\Xi$ be the true MV coordinates and $\hat{\Xi}=\Xi+\Delta$ be the interpolated coordinates. Under the Householder-stabilized system, the reconstruction error for the orthogonal projector $P=UU^{T}$ satisfies
@@
Let $\Psi(\Xi)=\widetilde{U}(\Xi)\widetilde{U}(\Xi)^{T}$ be the mapping from the coordinate domain to the Grassmannian. Define the line segment $\gamma(t)=\Xi+t\Delta$ for $t\in[0,1]$. By the Mean Value Theorem for matrix-valued functions, we have
@@
Taking the Frobenius norm on both sides yields
@@
To bound the differential norm at any point $X\in\{\gamma(t)\}$, we evaluate $d\Psi(X)[\Delta]=(d\widetilde{U})\widetilde{U}^{T}+\widetilde{U}(d\widetilde{U})^{T}$. Using the reconstruction formula $\widetilde{U}(X)=\begin{bmatrix}I_{p}\\ X\end{bmatrix}S(X)$ where $S(X)=(I_{p}+X^{T}X)^{-1/2}$, the differential is
@@
Its unique solution is given by the integral representation
@@
Taking the spectral norm and applying the bound $|e^{-S^{-1}\tau}|_{2}\leq e^{-\tau/|S|_{2}}$, the scalar integral can be evaluated to produce
@@
Substituting $d\tilde{U}$ into $d\Psi$, we decompose it into a principal part and a remainder
@@
Let $A=D_{1}\tilde{U}^{T}=\begin{bmatrix}0\\ \Delta S\end{bmatrix}\tilde{U}^{T}$. Applying the triangle inequality to the remainder $\mathcal{R}=D_{2}\tilde{U}^{T}+\tilde{U}D_{2}^{T}$ gives
@@
Thus the remainder is rigorously bounded by $\|\mathcal{R}\|_{F}=\mathcal{O}(\|X\|_{2}\|\Delta\|_{F})$. Next, we evaluate $\|A+A^{T}\|_{F}$. By the properties of the Frobenius norm,
@@
Thus, $\operatorname{tr}(A^{2})=\operatorname{tr}(SX^{T}\Delta S^{2}X^{T}\Delta S)$. By Cauchy-Schwarz inequality for trace inner products $|\operatorname{tr}(YZ)|\leq\|Y\|_{F}\|Z\|_{F}$, letting $Y=SX^{T}\Delta S$ and $Z=SX^{T}\Delta S$, we get
@@
Substituting back into (16), we obtain
@@
Using the sub-multiplicative property $\|\Delta S\|_{F}\leq\|S\|_{2}\|\Delta\|_{F}$,
@@
Combining the bounds for the principal part and the remainder $\mathcal{R}$, the differential norm at $X$ is bounded by
@@
Returning to the supremum in (14), since $X\in\{\Xi+t\Delta\}$, we have $\|X\|_{2}\leq\|\Xi\|_{2}+\|\Delta\|_{2}$. Furthermore, as established in Lemma 2, the matrix function $S(\cdot)$ is smooth, ensuring $\|S(X)\|_{2}=\|S(\Xi)\|_{2}+\mathcal{O}(\|\Delta\|_{F})$. Since the upper block of $\widetilde{U}(\Xi)$ is $S(\Xi)$, we have $\|\widetilde{U}_{1}\|_{2}=\|S(\Xi)\|_{2}$. Substituting these limits into the integral bound completes the strict derivation